How much positive and negative charge is there in a cup of water

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
Let us consider 1 cup = 250 mL = 250g
We know Density of Water = 1 \ \frac{g}{cc} (or) 1 \ \frac{g}{mL^3}
where g=grams, cc= centimeter cube, mL= mili litre
 
We know the formula of water is {{H}__2}O
Therefore in 18g of {{H}__2}O there are 8 protons and 8 electrons.
=> The number of protons and electrons each in 250 g of water are \frac{8\ast250}{18}\approx 111
=> There are 111 protons and 111 electrons in 250 mL of water
We know charge of 1 electron = -1.6\ast10^{-19} \ C and
charge of 1 proton =  +1.6\ast10^{-19} \ C
Therefore 
positive charge in a cup of water = 111\ast(+1.6\ast10^{-19})\approx+118\ast10^{-19}=+1.18*10^{-17} C
and
positive charge in cup of water = 111\ast(-1.6\ast10^{-19})\approx-118\ast10^{-19}=-1.18*10^{-17} C
 
Regards
Arun (askIITians forum expert)